Write the equation of the line that passes through (8,5) with a slope of -1 over 2

thank you

Answer by Alan3354(69443)

Write the equation of the line that passes through (8,5) with a slope of -1 over 2
----------
These are ALWAYS done like this:
Use y-y1 = m*(x-x1) m is the slope and (x1,y1) is the point
y-5 = (-1/2)*(x-8)
y = (-1/2)x + 9 (slope-intercept form)
-------------
x + 2y = 18 (standard form)
------------